DATA SHEET – EXPERIMENT 3A 00 Cu s Oo 03 062 08880 Simple Empirical Formula Determination for Magnesium Oxide Trial 1 1. MassOVORIO Class Data for Mg:0 Mole Ratios Class mean: Class standard deviation:4

PLEASE WRITE THE FORMULA!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
Answer #1

Empirical formula is the formula that gives the proportion (ratio) of the elements in the compound. It does not give the actual number of atoms of each element of the compound. It only gives the ratio. Therefore empirical formula may be different than molecular formula.

Now let us go for solving the problem.

On checking trial 1 data that has been enetered, there is an error in the moles of Mg. The value used for calculating the moles is 0.184 g, the correct value should be 0.105 g. The mass of Mg taken in the experiment is what was added in the beginning. After the reaction, the new mass of the product is 0.184 g.

So, the moles of Mg = mass of Mg/molar mass of Mg = 0.105g/24.31g/mol = 0.00432 mol

Moles of O = mass /molar mass = 0.079g/16 g/mol = 0.00494 mol

For finding the empirical formula, we divide the moles of each element by the element that has the least moles.

So we will divide all the moles of elements by the moles of Mg

For Mg = 0.00432/0.00432 = 1

For O = 0.00494/ 0.00432 = 1.14

Now we need to convert these numbers into nearest whole number.

Mor Mg it will be 1

For O we can convert 1.14 to 1 ( 1 is the nearest whole number to 1.14)

So the empirical formula is Mg1O1 = MgO

For trial 2

Moles of Mg = 0.108g/ 24.31 g/mol = 0.00444 mol

Moles of O = 0.069g/16g/mol = 0.00431 mol

Now in this case, moles of O are less. So we will divide by moles of O

For Mg 0.00444/0.00431 = 1.03 =1 (nearest whole number)

For O 0.00431/0.00431 = 1

So the empirical formula = MgO

Note: the question also asks for the mole ratio of Mg:O. For this, the moles of Mg should be divided by moles of O in both trials.

Mole ratio for trial 1

Moles of Mg/moles of O = 0.00432/0.00494 = 0.874

For trial 2

= O.00444/ 0.00431 = 1.03

Average mole ratio = 0.874+1.03/ 2 = 0.95

Nearest whole number = 1

Empirical formula = MgO

Part2

Trial 1

Mass of CuSO4.xH2O = 44.954g - 43.920g = 1.034g

Mass of water lost = 44.954 - 44.551 = 0.403g

Mass of CuSO4 = 1.034 - 0.403 = 0.631 g

Molar mass of CuSO4 = 159.61g/mol

Moles of CuSO4 = 0.631 g/ 159.61 g/mol = 0.00395 mol

Moles of water = 0.403g / 18g/mol = 0.0224 mol

For finding the # of moles of water of hydration, divide the moles of water by moles of CuSO4 = 0.224/ 0.00395 = 5.67

For trial 2

Mass of CuSO4.xH2O = 44.821- 43.933 = 0.888 g

Mass of water lost = 44.821 - 44.725 = 0.096 g

Mass of CuSO4 = 0.888 - 0.096 = 0.792 g

Moles of CuSO4 = 0.792 g/ 159.61 g/mol = 0.00496 mol

Moles of water = 0.096g/18g/mol = 0.0053 mol

# of moles of water of hydration = 1.075

There seems to be a large difference between the value in trial 1 and trial 2. The few sources of error can be

1. The mass measurement was incorrect

2. The source of CuSO4.xH2O in trial 2 was different than that in trial 1

3. The crucible was not heated properly, so the water loss was not complete.

Please recheck all the data before reporting the results